What is the final temperature of a drink if the initial 1.30 kg of water is at a temperature of 25.0 degC and 0.0800 kg of ice at 0 degC is placed in it?
Answer #1

Heat Gained by Ice = Heat Lost by water

miceLf+Micecw(T-0) = mwcw(To-T)

0.08*(3.34*105) +0.08*4186*T = 1.3*4186*(25-T)

5776.68T = 109325

T=18.925oC